# Prismatic-Bending Transformable (PBT) Joint for a Modular, Foldable Manipulator with Enhanced Reachability and Dexterity

### v1

@@ Abstract
Robotic manipulators, traditionally designed with classical joint-link articulated structures, excel in industrial applications but face challenges in human-centered and general-purpose tasks requiring greater dexterity and adaptability. Addressing these limitations, we introduce the Prismatic-Bending Transformable (PBT) Joint, a novel design inspired by the scissors mechanism, enabling transformable kinematic chains. Each PBT joint module provides three degrees of freedom—bending, rotation, and elongation/contraction—allowing scalable and reconfigurable assemblies to form diverse kinematic configurations tailored to specific tasks. This innovative design surpasses conventional systems, delivering superior flexibility and performance across various applications. We present the design, modeling, and experimental validation of the PBT joint, demonstrating its integration into modular and foldable robotic arms. The PBT joint functions as a single SKU, enabling manipulators to be constructed entirely from standardized PBT joints without additional customized components. It also serves as a modular extension for existing systems, such as wrist modules, streamlining design, deployment, transportation, and maintenance. Three sizes—large, medium, and small—have been developed and integrated into robotic manipulators, highlighting their enhanced dexterity, reachability, and adaptability for manipulation tasks. This work represents a significant advancement in robotic design, offering scalable and efficient solutions for dynamic and unstructured environments.
@@ I Introduction
ROBOTIC manipulators are indispensable tools that empower robots to interact with the physical world, much like human arms facilitate our own interactions. Traditional robotic manipulators are built on two fundamental joint types—prismatic and revolute—connected by rigid links to form desired kinematic chains [1]. These systems have resulted in six widely recognized geometric configurations: articulated, spherical, SCARA, cylindrical, Cartesian, and parallel manipulators [2]. Among these, articulated manipulators are particularly prevalent due to their intuitive kinematics, making them well-suited for high-accuracy, high-payload tasks [3, 4]. However, human-centered environments, such as domestic or service applications, present unique challenges that prioritize adaptability, dexterity, obstacle avoidance, and reachability over traditional metrics such as payload capacity and precision [5, 6, 7, 8]. Additionally, while robots performing tasks in structured environments primarily focus on avoiding obstacles at the end-effector, in unstructured environments, the body configuration of the manipulator is significantly influenced by the complexity and dynamics of the environment. Achieving greater flexibility in body configuration while maintaining terminal manipulability is crucial for enabling robots to effectively perform tasks in the unstructured settings [2, 5, 9].
@@
To enhance the dexterity and reachability of robotic manipulators in unstructured environments and human-centered tasks, researchers have explored two distinct solutions: continuum manipulators and foldable robots, both offering innovative alternatives to traditional articulated manipulators. Continuum manipulators, inspired by biological structures such as elephant trunks and octopus arms, achieve smooth, adaptive motion through continuously deformable materials or mechanisms [10, 11]. These systems provide multi-directional bending, enabling exceptional obstacle avoidance and adaptability in unstructured environments [12, 13]. Our previous work has also demonstrated the potential of continuum robots to enhance dexterity in confined spaces. For example, a soft wrist design utilizing tendon differential actuation enables human wrist-like dexterity and manipulability for teleoperated swab sampling [14, 15]. By incorporating a tendon-jamming mechanism, the stiffness of the continuum manipulator becomes adjustable while maintaining excellent motion dexterity for operations in confined spaces [16].
Foldable manipulators leverage the unique capabilities of foldable mechanisms to enhance the reachability and adaptability of robotic systems [17]. By allowing their structures to extend, fold, or reshape, these manipulators offer exceptional versatility and can be tailored to meet task-specific requirements. The principles of modularity and foldability enable quick deployment, efficient space-saving storage, and effective obstacle avoidance, making them ideal for use in dynamic and constrained environments [18]. Foldable manipulators have demonstrated their unique advantages in various applications. For instance, a foldable manipulator integrated with UAVs showcases its ability to extend into confined spaces for tasks such as object retrieval and inspections, highlighting its potential in challenging scenarios like search and rescue operations [19]. Furthermore, foldable manipulators can effectively merge locomotion and manipulation functions, making them a promising solution for space exploration. By combining these capabilities, such manipulators can navigate and interact with complex extraterrestrial environments, as demonstrated in ReachBot’s integration for extraterrestrial missions [20].
Although continuum and foldable manipulators each offer distinct advantages, neither approach successfully integrates multi-directional bending with foldability into a single system—a promising avenue for significantly enhancing the dexterity and reachability of robotic manipulators.
To address this need, this paper introduces the Prismatic-Bending Transformable (PBT) Joint, a novel mechanism that combines multi-directional bending flexibility with foldability inspired by the scissors mechanism. Each PBT Joint provides three degrees of freedom—bending, rotation, and elongation/contraction—with transformable prismatic and bending modes. This design enables the assembly of modular and foldable robotic manipulators that are highly adaptable to unstructured environments and diverse task requirements. Details of the design, modeling, and experimental validation of the Prismatic-Bending Transformable (PBT) Joint are discussed, showcasing its ability to integrate seamlessly into robotic arms while significantly enhancing their dexterity, manipulability, and obstacle avoidance. Through quantitative analysis of its dexterity and manipulability, the PBT Joint integrated manipulator demonstrates superior performance in adapting to diverse tasks and complex environments. Serving as a single standardized SKU, the PBT Joint enables the construction of complete manipulators without requiring additional customized components. It also functions as a modular extension for existing robotic systems, such as a PBT wrist module, further broadening its applicability. To meet various application requirements, three sizes of PBT Joints—large, medium, and small—have been developed and tested in different robotic configurations. These implementations underscore the joint’s adaptability, reconfigurability, and effectiveness, positioning it as a promising solution for general-purpose robotic tasks and dynamic, human-centered applications.
@@
  • Novel Joint Design: Development of the Prismatic-Bending Transformable (PBT) Joint, a modular unit integrating multi-directional bending, rotation, and elongation/contraction with a 3D direction maintenance mechanism, offering adaptable dexterity and task-specific customization.
  • Enhanced Robotic Manipulator Architecture: Development of a modular and foldable robotic manipulator architecture using PBT joints, enabling enhanced dexterity, manipulability, and obstacle avoidance, with applications as an extension wrist or a foldable two-joint manipulator for unstructured environments.
  • Detailed analysis, modeling, planning, and experimental validation demonstrating the feasibility, performance, and versatility of the PBT joint in enabling robotic arms to perform a wide range of tasks with enhanced dexterity and manipulability.
The paper is organized as follows: Section II introduces the transformation between prismatic and bending modes in the PBT design. Section III details the PBT Joint’s design and its integration into modular manipulators. Section IV analyzes the kinematics and manipulability of a dual-unit system. Section V presents experimental validations for single joints and modular systems, followed by the conclusion and future work.
@@ II Concept of PBT Joint
This section introduces the motion and degrees of freedom (DOFs) of a single PBT joint and its modular extension. As illustrated in Figure 2A, the proposed PBT joint employs a scissors-inspired mechanism to achieve reliable prismatic and bending motion, providing two correlated DOFs [21, 22, 23, 24]. This mechanism forms the basic PBT joint, enabling prismatic and bending transformations. Self-rotation is achieved by integrating a perpendicular rotation actuator with the core scissors mechanism [25].
@@ III Design of PBT Joint and Modular Manipulator Architecture
Figure 4: Structure and Assembly of the Two-Module PBT Joint Modular Manipulator (A) Partial sectional front view of the modular manipulator: 1 represents the timing belt, 2 is the servo motor driving the linear motion of the PBT joint, and 3 represents the driving gear of the PBT joint directly connected to the servo motor. (B) Full view of the modular manipulator: 4 represents the servo motors driving the revolute motion of the PBT joint, 5 represents the servo motor driving the revolute motion of the medium PBT joint, and 6 represents the servo motor driving the linear motion of the medium PBT joint. (C) Partial sectional view of the modular manipulator: The simplified link and joint structure are illustrated for the large PBT joint. (D) Partial exploded sectional view of the modular manipulator: From bottom to top are the large revolute module, large PBT module, medium revolute module, and medium PBT module. (E) Two potential structures for folding are shown. Force singularity occurs in the structure in the second row.
@@ IV-A Inverse Kinematics
Figure 5: Reachability and Manipulability Analysis (A) Reachability set of a single PBT joint: (i) and (ii) correspond to the prismatic and bending modes, respectively. (B) Single PBT joint (wrist) Mounted on a Collaborative Robot Arm: The reachability set in a compact environment is extended. (i) visualizes the environment, while (ii) and (iii) show workspaces with and without the mounted PBT joint. Scatter plots represent end-effector positions from 10,000 uniformly sampled joint configurations, excluding those with collisions. In (ii), sampling density is increased in the cavity area to better capture reachability. (C) Translational Manipulability Distribution in Task Space: (i)-(iv) correspond to the PP, PB, BP, and BB modes defined in Fig. 2. Translational manipulability quantifies the modular manipulator’s ability to move the end-effector in different directions in the task space, calculated using (5). (D) Workspace in a Compact Environment: (i) visualizes the environment. (ii) and (iii) show the BP and BB mode workspaces, represented by scatter plots from a top-down view. The scatter plots are generated from 30,000 randomly sampled joint configurations, excluding collisions. (E) Workspace in a Clustered Environment: (i) visualizes the environment. (ii) and (iii) show the BP and BB mode workspaces, represented by scatter plots from a side view. Scatter plots are based on 10,000 uniformly sampled joint configurations, excluding collisions.
@@ IV-B Reachability and Manipulability Analysis
Fig. 5A shows the reachable set of a single PBT joint, from which we can see that the P mode has a reachable set of a line and the B mode has sphere reachable set. Fig. 5B demonstrates the capability of PBT joints to extend a robot’s workspace in a compact environment. The PBT joint allows manipulation in a cavity area whose opening is opposite to the robot’s base. The robot end-effector is unable to enter the cavity area without the assistance of a PBT joint.
Fig. 5C visualizes the translational manipulability of distributed in the end effector’s reachable positions. Translational manipulability describes the required effort to drive the robot towards all possible task-space directions at a given configuration. It is quantified here following the definition in [28]
@@
The advantages of mode transformation becomes evident in the presence of obstacles. Prismatic mode allows PBT joints to shrink its dimensions, enhancing flexibility in compact environments, while bending mode enables the end-effector to navigate around obstacles, offering greater dexterity in clustered scenarios. Since no general quantitative measure for collision-free dexterity exists, a qualitative analysis is provided.
Figure 5D illustrates performance in a narrow tunnel. The BB mode results in a split, disconnected workspace, limiting the end-effector’s ability to transition between zones. In contrast, the BP mode creates a connected workspace, demonstrating greater flexibility in confined spaces. Figure 5E highlights performance in clustered environments. The BP mode cannot reach beyond an obstacle, whereas the BB mode achieves a significant collision-free workspace on the other side, showcasing the bending mode’s dexterity for navigating obstacles in cluttered scenarios.
@@ V-A Demonstration of The One Module PBT Wrist
The PBT wrist is a small-sized PBT joint (Fig. 3). It also uses the same link configuration (Fig. 4E) to avoid singularities in force during lifting and lowering motions, with its parameters as shown in Tab. I. Unlike the previously shown PBT joint, the wrist’s revolute joint is not driven by a synchronous belt but rather by directly mounting a Dynamixel servo motor (M226) on the revolute axis, which directly drives the revolute motion [29]. Thus, the reduction ratio of the revolute joint is 1. Another identical Dynamixel servo motor is installed on the orientation-maintaining mechanism, driving the wrist’s linear motion through a gear structure with a reduction ratio of 2. Both motors are connected to the PC via a servo driver board, enabling simple control (Fig. 1B and 1C) [30].

### v2

@@ Abstract
Robotic manipulators, traditionally designed with classical joint-link articulated structures, excel in industrial applications but face challenges in human-centered and general-purpose tasks requiring greater dexterity and adaptability. To address these challenges, we propose the Prismatic-Bending Transformable (PBT) Joint—a novel, scissors-inspired mechanism with directional maintenance capability that provides bending, rotation, and elongation/contraction within a single module. This design enables transformable kinematic chains that are modular, reconfigurable, and scalable for diverse tasks. We detail the mechanical design, optimization, kinematic and dynamic modeling, and experimental validation of the PBT joint, demonstrating its integration into foldable, modular robotic manipulators. The PBT joint functions as a single stock keeping unit (SKU), enabling manipulators to be constructed entirely from standardized PBT joints. It also serves as a modular extension for existing systems, such as wrist modules, streamlining design, deployment, transportation, and maintenance. Three joint sizes have been developed and tested, showcasing enhanced dexterity, reachability, and adaptability, particularly in confined and cluttered spaces. This work presents a promising approach to robotic manipulator development, providing a compact and versatile solution for operation in dynamic and constrained environments.
@@ I Introduction
Robotic manipulators are essential tools that enable robots to interact with the physical world, much like human arms facilitate our own interactions. Traditional manipulators are typically constructed using two fundamental joint types—prismatic and revolute—connected by rigid links to form various kinematic chains [1]. These have led to the development of six well-known geometric configurations: articulated, spherical, SCARA, cylindrical, Cartesian, and parallel manipulators [2]. Among these, articulated manipulators are the most widely adopted due to their intuitive kinematics and suitability for high-accuracy, high-payload tasks [3, 4].
However, the increasing demand for robots in human-centered environments, such as domestic, assistive, or service applications, introduces new challenges that emphasize adaptability, dexterity, obstacle avoidance, and reachability over traditional metrics like precision and payload capacity [5, 6, 7, 8]. Moreover, while robots operating in structured settings can often focus on avoiding obstacles only at the end-effector, those deployed in unstructured environments must consider the full-body configuration of the manipulator in response to dynamic surroundings. Achieving greater flexibility in body posture while maintaining end-effector manipulability is therefore essential for effective manipulation [2, 5, 9].
@@
To address the challenges of dexterity and adaptability in unstructured and human-centered environments, researchers have primarily explored two innovative design paradigms: continuum manipulators and foldable robotic mechanisms. Continuum manipulators, inspired by biological systems such as elephant trunks and octopus arms, enable smooth and adaptive motion through continuously deformable bodies or tendon-driven structures [10, 11]. These systems provide inherent compliance and multi-directional bending, making them especially well-suited for navigating complex, cluttered environments [12, 13]. Recent developments in this field have focused on enhancing modeling, control strategies, and task-specific designs for manipulation in constrained and irregular spaces [12]. Our previous work has contributed to this area through the development of a tendon-driven soft wrist with human-like articulation for teleoperated nasopharyngeal swab sampling [14, 15]. We also introduced a tendon-jamming mechanism to achieve tunable stiffness, allowing the structure to adapt to varying load conditions while maintaining dexterous motion [16].
In parallel, foldable manipulators have emerged as a compelling solution for achieving reconfigurability, compact deployment, and structural adaptability [17]. These systems utilize mechanisms capable of extending, folding, or reshaping to optimize workspace usage and enhance maneuverability in constrained environments. They are particularly valuable in domains such as disaster response, mobile robotics, and space exploration [18]. Recent developments have highlighted innovative applications of origami-inspired designs—for example, Kim et al. proposed a self-locking robotic arm that folds flat and deploys into a functional manipulator, achieving both portability and structural rigidity [19]. Similarly, Chen et al. introduced ReachBot, a robot designed for extraterrestrial missions that integrates locomotion and manipulation through deployable booms, enabling interaction with distant or vertical surfaces [20]. Suthar and Jung developed a foldable robot arm for drones using a twisted string actuator, demonstrating a lightweight and compact solution well-suited for aerial manipulation tasks [21]. Complementary to these approaches, Dai et al. have advanced the field of metamorphic and reconfigurable mechanisms, exploring SLPM-based deployable robots and modular systems capable of changing their kinematic structure to meet varying task demands [22, 23].
Despite these advancements, most existing designs fall short in simultaneously providing post-deployment dexterity and scalable modularity. Continuum systems often face challenges in precise control and complex actuation schemes, while foldable systems typically lack adaptability once deployed. Moreover, few platforms successfully integrate multi-directional bending with structural transformability in a single unified architecture—an integration that could significantly enhance the versatility and functional performance of robotic manipulators operating in dynamic and unstructured environments.
To bridge this gap, we propose the Prismatic-Bending Transformable (PBT) Joint—a novel, scissor-inspired mechanism that combines the multi-directional flexibility and foldability. The joint supports bending, rotational, and prismatic motions, enabling smooth transitions between prismatic and bending configurations. This architecture facilitates compact, reconfigurable manipulators that expand into highly dexterous structures when deployed. We present the joint’s design, along with its kinematic and dynamic modeling, and validate its performance through experimental evaluation. Quantitative analysis and comparisons demonstrates improvements in dexterity, manipulability, and workspace reach. Designed as a standardized SKU, the PBT Joint enables scalable manipulator construction without the need for custom components, and it can also function as a modular extension, such as a reconfigurable wrist. We developed and tested three joint sizes—large, medium, and small—across multiple configurations, confirming the joint’s scalability, versatility, and effectiveness for general-purpose manipulation in dynamic, human-centered environments.
@@
  • Development of the Prismatic-Bending Transformable (PBT) Joint, a modular unit integrating multi-directional bending, rotation, and elongation/contraction with a 3D direction maintenance mechanism, offering adaptable dexterity and task-specific customization.
  • Development of a modular and foldable robotic manipulator architecture using PBT joints, enabling enhanced dexterity, manipulability, and obstacle avoidance, with applications as an extension wrist or a foldable two-joint manipulator for unstructured environments.
  • Comprehensive analysis, including kinematic, manipulability, workspace, singularity avoidance and dynamic modeling, guides PBT joint and its enabled manipulator design.
The paper is organized as follows: Section II introduces the concept of the PBT Joint. Section III describes the joint’s design and its integration into modular manipulators. Section IV analyzes the kinematics, dynamics, and manipulability. Section V presents experimental validations for both single joints and modular systems. The paper concludes with a summary and discussion of future work.
@@ II Concept of PBT Joint
This section introduces the motion and degrees of freedom (DOFs) of a single PBT joint and its modular extension. As illustrated in Figure 2A, the proposed PBT joint employs a scissors-inspired mechanism to achieve reliable prismatic and bending motion, providing two correlated DOFs [24, 25, 26]. This mechanism forms the basic PBT joint, enabling prismatic and bending transformations. Self-rotation is achieved by integrating a perpendicular rotation actuator with the core scissors mechanism [27].
@@ III Design of PBT Joint and Modular Manipulator Architecture
Figure 4: Structure and Assembly of the Two-Module PBT Joint Modular Manipulator (A) Partial sectional front view of the modular manipulator: 1 represents the timing belt, 2 is the servo motor driving the linear motion of the PBT joint, and 3 represents the driving gear of the PBT joint directly connected to the servo motor. (B) Partial exploded sectional view of the modular manipulator: From bottom to top are the large revolute module, large PBT module, medium revolute module, and medium PBT module. Two potential structures for folding are shown. Force singularity occurs in the structure in the second row. (C) Full view of the modular manipulator: 4 represents the servo motors driving the revolute motion of the PBT joint, 5 represents the servo motor driving the revolute motion of the medium PBT joint, and 6 represents the servo motor driving the linear motion of the medium PBT joint.
@@ IV-A Inverse Kinematics
Figure 5: Reachability and Manipulability Analysis of PBT Joints (A) Reachability set of a single PBT joint: (i) and (ii) correspond to the prismatic and bending modes, respectively. (B) Translational Manipulability Distribution in Task Space of dual-unit PBT: (i)-(iv) correspond to the PP, PB, BP, and BB modes defined in Fig. 2. Comparing different motion modes with a conventional UR5 arm in an compact (C) and a clustered (D) environments. (i) Environment visualization, the dual-unit PBT and UR5 are scaled to a normalized stretched length. (ii)-(iV) show the BP mode, BB mode and UR5 workspaces. The scatter plots are generated from 30,000 randomly sampled joint configurations, excluding collisions. Extended conventional robot’s workspace with PBT units (E) The reachability set in a compact environment is extended with a PBT joint mounted on a UR5. (i) visualizes the environment, while (ii) and (iii) show workspaces with and without the mounted PBT joint. In (ii), sampling density is increased in the cavity area to better capture reachability.
@@ IV-B Reachability and Manipulability Analysis
Fig. 5A shows the reachable set of a single PBT joint, from which we can see that the P mode has a reachable set of a line and the B mode has sphere reachable set. Fig. 5B visualizes the translational manipulability of a dual-unit PBT distributed in the end effector’s reachable positions. Translational manipulability describes the required effort to drive the robot towards all possible task-space directions at a given configuration. It is quantified here following the definition in [30]
@@
The advantages of mode transformation becomes evident in the presence of obstacles. Prismatic mode allows PBT joints to shrink its dimensions, enhancing flexibility in compact environments, while bending mode enables the end-effector to navigate around obstacles, offering greater dexterity in clustered scenarios. Since no general quantitative measure for collision-free dexterity exists, a qualitative analysis is provided. Figure 5C shows performance in a narrow tunnel. BB mode creates a split workspace, limiting transitions, while BP mode maintains connectivity for better flexibility. The UR5, despite a similar size, twists excessively and has the most fragmented workspace. Figure 5D illustrates performance in clustered environments. BP mode cannot reach beyond obstacles, while BB mode achieves a larger collision-free workspace, demonstrating superior dexterity. The UR5 has a smaller reach on the other side comparing to the BB mode.
Fig. 5E demonstrates the capability of PBT joints to extend a robot’s workspace in a compact environment. The PBT joint allows manipulation in a cavity area whose opening is opposite to the robot’s base. The robot end-effector is unable to enter the cavity area without the assistance of a PBT joint.
IV-C Inverse Dynamics
The inverse dynamics refers to the problem of determining the joint torques $\tau$ given the robot motion $\theta,\dot{\theta},\ddot{\theta}$. The recursive Newton-Euler algorithm is efficient to solve the serial robot manipulators. As the bending mode is similer to serial manipulator, we focus on the dynamics of prismatic mode of PBT joint.
As the PBT joint has a symmetrical structure, we use Lagrangian method
$\frac{d}{dt}\left(\frac{\partial L}{\partial\dot{\theta}}\right)-\frac{\partial L}{\partial\theta}=\tau$ (6)
to obtain its dynamics.
Assuming ${}^{\rm w}v_{i-1},^{\rm w}\omega_{i-1}$ represent the linear and angular velocity from the previous link in the world frame, the velocity of the center of mass and angular velocity of the PBT joint is
$\begin{cases}{}^{\rm w}v_{i}=^{\rm w}v_{i-1}+^{i-1}R_{i}\begin{bmatrix}0\\ 0\\ -lsin\frac{\theta}{2}\dot{\theta}\end{bmatrix}\\ {}^{\rm w}\omega_{i}=^{\rm w}\omega_{i-1}\end{cases}$ (7)
where ${}^{i-1}R_{i}$ represents the rotation matrix from frame $i-1$ to the PBT frame $i$. The kinetic and potential energy of the PBT joint is
$\begin{cases}K_{i}=\frac{1}{2}m_{i}v_{i}^{T}v_{i}+\frac{1}{2}\omega_{i}^{T}I_{i}\omega_{i}\\ U_{i}=m_{i}g^{T}r_{i}\end{cases}$ (8)
where $I_{i}=\text{diag}(\frac{1}{2}m_{i}l^{2},\frac{1}{2}m_{i}l^{2}\cos^{2}\frac{\theta}{2},\frac{1}{2}m_{i}l^{2}\sin^{2}\frac{\theta}{2})$ is the inertial matrix. It is straightforward to obtain the Lagrangian operator on the PBT joint as
$\tau=$
$m_{i}l^{2}\sin^{2}\frac{\theta}{2}\ddot{\theta}+m_{i}l^{2}\sin\frac{\theta}{2}\cos\frac{\theta}{2}\dot{\theta}^{2}-\frac{1}{2}m_{i}l^{2}\sin^{2}\frac{\theta}{2}\dot{\theta}^{2}$
$+m_{i}(^{\rm pbt}\dot{v}_{i-1})^{T}\begin{bmatrix}0\\ 0\\ -l\sin\frac{\theta}{2}\end{bmatrix}-\frac{1}{2}m_{i}lg^{T}\cdot^{i-1}R_{i}\begin{bmatrix}0\\ 0\\ \sin\frac{\theta}{2}\end{bmatrix}$ (9)
$-\frac{1}{2}\omega_{i-1}^{T}\frac{\partial I_{i}}{\partial\theta}\omega_{i}+m_{i}(^{\rm pbt}v_{i-1})^{T}\begin{bmatrix}0\\ 0\\ \frac{\dot{\theta}l}{2}(\sin\frac{\theta}{2}-\cos\frac{\theta}{2})\end{bmatrix}.$
where the ${}^{\rm pbt}\dot{v}_{i-1}$ represents the velocity of previous link in the PBT joint’s frame. It’s obvious that when the base is static and $\theta=\pi$, the maximum required torque to lift the PBT joint is $\frac{1}{2}m_{i}gl$. As analyzed in Fig. 4B, if the applied torque is insufficient, lifting the joint may be impossible in certain kinematic configurations.
@@ V-A Demonstration of The One Module PBT Wrist
The PBT wrist is a small-sized PBT joint (Fig. 3). It also uses the same link configuration (Fig. 4E) to avoid singularities in force during lifting and lowering motions, with its parameters as shown in Tab. I. Unlike the previously shown PBT joint, the wrist’s revolute joint is not driven by a synchronous belt but rather by directly mounting a Dynamixel servo motor (M226) on the revolute axis, which directly drives the revolute motion. Thus, the reduction ratio of the revolute joint is 1. Another identical Dynamixel servo motor is installed on the orientation-maintaining mechanism, driving the wrist’s linear motion through a gear structure with a reduction ratio of 2. Both motors are connected to the PC via a servo driver board, enabling simple control (Fig. 1B and 1C) [31].